OpenStudy (anonymous):

A ball is thrown horizontally at a speed of 24 meters per second from the top of a cliff. If the ball hits the ground 4.0 seconds later, approximately how high is the cliff?

OpenStudy (abhisar):

Hello @deimcc ! Welcome to OpenStudy !

OpenStudy (abhisar):

Using the equation \(\sf s=ut+0.5at^2\) \(\sf s=0\times 4+0.5\times 9.8\times 16\\ \Rightarrow s=78.4 m\)